The most recent financial statements for Moose Tours, Inc., appear below. Sales for 2016 are projected to grow by 20 percent. Interest expense will remain constant; the tax rate and the dividend payout rate will also remain constant. Costs, other expenses, current assets, fixed assets, and accounts payable increase spontaneously with sales. MOOSE TOURS, INC. 2015 Income Statement Sales Costs Other expenses $751,000 586,000 22,000 Earnings before interest and taxes Interest expense $143,000 18,000 Taxable income Taxes (40%) $125,000 50,000 Net income $75,000 Dividends Addition to retained earnings $30,000 45,000 MOOSE TOURS, INC Balance Sheet as of December 31, 2015 Assets Liabilities and Owners Equity Current assets Current liabilities Cash $ 21,040 Accounts payable 55,200 14,400 69,600 S 134,000 33,360 Notes payable 70,320 Inventory Total Total $124,720 Long-term debt Owners equity Common stock and paid-in surplus Retained earnings Fixed assets Net plant and equipment $240,000 $120,000 41,120 S161,120 S364,720 Total assets $364,720 Total liabilities and owners equity

If the firm is operating at full capacity and no new debt or equity is issued, what external financing is needed to support the 20 percent growth rate in sales? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

  External financing needed $   
Answer #1

EFN = Increase in assets – Increase in liabilities – Increase in retained earnings

EFN = 72944-11040- 55296 = 6608

WORKINGS

Step 1:

Increase in retained earnings

Sales = 751000*120%=901200

Less: costs 586000*120%=703200

Other expenses = 22000*120%=26400

EBIT = 171600

Less: Interest 18000

Taxable income =153600

Tax = 40% = 61440

Net Income = 92160

Dividend = (30000*92160/75000)= 36864

Addition to retained earnings = 55296

Step 2: Increase in assets = (Current + Fixed assets)*20%

= (124720+240000)*20%

=72944

Increase in spontaneous liabilities = Accounts payable *20%

=55200*20%

=11040

Step 3:

EFN = Increase in assets – Increase in liabilities – Increase in retained earnings

EFN = 72944-11040- 55296 = 6608